# On the Partitioning of GPU Power among Multi-Instances

### v1

@@ Abstract
Efficient power management in cloud data centers is essential for reducing costs, enhancing performance, and minimizing environmental impact. GPUs, critical for tasks like machine learning (ML) and GenAI, are major contributors to power consumption. NVIDIA’s Multi-Instance GPU (MIG) technology improves GPU utilization by enabling isolated partitions with per-partition resource tracking, facilitating GPU sharing by multiple tenants. However, accurately apportioning GPU power consumption among MIG instances remains challenging due to a lack of hardware support. This paper addresses this challenge by developing software methods to estimate power usage per MIG partition. We analyze NVIDIA GPU utilization metrics and find that light-weight methods with good accuracy can be difficult to construct. We hence explore the use of ML-based power models to enable accurate, partition-level power estimation. Our findings reveal that a single generic offline power model or modeling method is not applicable across diverse workloads, especially with concurrent MIG usage, and that online models constructed using partition-level utilization metrics of workloads under execution can significantly improve accuracy. Using NVIDIA A100 GPUs, we demonstrate this approach for accurate partition-level power estimation for workloads including matrix multiplication and Large Language Model inference, contributing to transparent and fair carbon reporting.
@@ I Introduction
Cloud data centers have become the backbone of modern computing, enabling a wide spectrum of applications and services, from social media platforms to complex scientific computations in a cost-efficient manner[1, 2]. As these centers grow in scale and complexity, efficient power management is gaining significant traction to control the rising energy costs and comply with mandatory carbon emission standards being introduced by governments across the world [3, 4]. Effective power management strategies are essential not only for reducing operational costs but also for enhancing system performance and minimizing the environmental impact due to them. Among the various components of a data center, Graphics Processing Units (GPUs) play a critical role, particularly in high-performance computing tasks such as machine learning, data analytics, and scientific simulations. With the increasing adoption of GenAI-based applications, the projected growth for data center GPU’s during 2023-28 is 34.6%[5], necessitating the development of efficient strategies for improving their utilization levels and energy efficiency.
@@
We next consider increasingly complex methods for our task at hand. Our first approach is to construct a universal power model that is applicable for a variety of common workloads and use it for independently estimating each partition’s power. It turns out that with ML methods such as Extreme Gradient Boosting (XGBoost)[11], errors can be high with unified models constructed using data sets from dissimilar multiple workloads, in non-MIG, full GPU mode. However, for unified models constructed from similar workloads, the error is substantially reduced. This highlights the potential advantages of applying workload-aware models, which we examine further. Interestingly, we demonstrate that individual workload-aware models do not significantly reduce prediction error when applied to MIG power prediction, whilst a unified model adds to predictivity. To that end, the availability of runtime workload monitored metrics to facilitate workload-aware model construction presents a compounding predicament to creating an accurate model that we highlight through this study. Finally, as a measure of improving accuracy, with all the methods, we scale the model-estimated MIG power values using aggregate measured GPU power, when available, and evaluate its efficacy.
@@
  • Design of approaches to partition GPU power among MIG instances & their evaluation, highlighting the challenges posed by the problem, and showing that lightweight approaches that can be implemented in a low-level module may not be possible.
@@ II Overview of NVIDIA GPU Architecture
Graphics Processing Units (GPUs), originally intended for accelerating computer graphics[12], have emerged as a cornerstone of high-performance computing due to their highly-parallel processor architecture, consisting of hundreds to thousands of small cores. In NVIDIA GPUs [^1] Unless otherwise specified, GPU henceforth refers to NVIDIA’s GPUs., the processing cores are organized into Streaming Multiprocessors (SMs)[13, 14, 10, 15]. Each SM consists of two types of cores, CUDA[16] and Tensor cores[17, 18], to support arithmetic operations of different types and precisions, such as 16-, 32-, and 64-bit floating points (FP16, FP32, and FP64 resp.) and 8- and 16-bit integers (INT8 and INT16, resp.). Each CUDA core is a general-purpose core that performs scalar operations per clock cycle, while tensor cores, designed to accelerate AI workloads, can perform mixed-precision matrix operations in a single cycle. Because of their simpler architecture, the CUDA cores significantly outnumber tensor cores in SM’s. GPUs also incorporate sophisticated memory hierarchies, including on-chip L1/L2 caches and shared memory, and large memory bandwidths for rapid data transfers between off-chip memory and on-chip caches. These features further improve performance in data-intensive tasks.
The above hardware support for parallelism is accompanied by a matching parallel software programming and execution model based on warps. A warp is a group of threads, typically 32, each of which is assigned to a distinct SM core and executes the same instruction sequence concurrently with other threads but on different data. The overall workload is broken down into warps, which are assigned to SMs. Each SM schedules the threads of a warp on its cores and executes them. Execution latency of warps is minimized by switching to a different ready warp if the executing wrap stalls. Such fine-grained multithreading allows the GPU to keep its cores fully utilized, even when certain threads experience delays. This single instruction multiple thread / data (SIMT / SIMD) model allows GPUs to leverage the inherent data-level parallelism present in many computational workloads, including AI.
Measuring and understanding the performance of a GPU requires monitoring its operation and use via various hardware metrics. GPU metrics include core clock speed (SMCLK), memory utilization (DRAMA), utilization of CUDA (FP64A, FP32A, FP16A, INT16A, etc.) & tensor compute cores (TENSO) and power consumption (POWER)[19]. These metrics provide insights into how efficiently the GPU is operating under different workloads and can be used to optimize performance and energy usage. In a data center, these GPU metrics can be accessed with minimal overhead via the NVIDIA Data Center GPU Manager (DCGM)[20].
@@
In contrast, NVIDIA Nsight Compute[21] is a more granular profiling tool that provides detailed, kernel-level analysis of GPU workloads. It allows developers to dive deep into performance metrics, such as memory access patterns, warp execution efficiency, and thread divergence, to extract a highly detailed view of the GPU’s inner workings. While Nsight Compute provides valuable insights for optimizing specific workloads, it incurs a higher computational overhead and is typically used in development and tuning rather than real-time monitoring. In this study, both DCGM and Nsight metrics are evaluated for their effectiveness in predicting power consumption, with a focus on balancing detail with efficiency for cloud data center applications.
Multi-Instance GPU (MIG): Recent generations of data center GPUs offer compute throughput of the order of a few hundred to a thousand TFLOPS[22, NVIDIA-H100-tput], which all workloads may not be capable of utilizing. To improve the overall utilization, extend access to wider set of users, and also reduce cost, different sharing modes namely, time-slicing, multi-process service (MPS)[23], and multi-instance GPU (MIG)[7] are supported by NVIDIA. Of these, only MIG offers complete performance and memory isolation. Under MIG, at the base level, a GPU’s SM’s are divided into seven and memory into eight equally-sized slices. At the next level, the base slices are combined in certain permissible counts to form larger partitions, referred to as GPU instances or simply partitions. The configurations, referred to as MIG profiles, permissible for the MIG partitions in an A100 with 80 GB memory are provided in Table I. Currently, a GPU can be divided into up to four partitions at a time. As discussed in I, when operating in MIG mode, GPU utilization metrics are reported at the partition level, while power consumption is available only for the full GPU. The rest of the paper discusses methods for distributing the aggregate power among the partitions.
@@ III GPU Power-Performance Characterization Studies and Modeling
In this section, we examine the various factors that affect power consumption in GPUs using NVIDIA A100 and V100 systems. We use different types of workloads such as various CUDA matrix multiplication kernels (MATMUL)[24], a TENSOR workload based on GPU Burn benchmark (GPUBurn)[25], and Large Language Model (LLM) inference systems[26]. MATMUL includes the evaluation of various matrix multiplication kernels (Kernels 1 – 10) across a broad spectrum of matrix sizes, ranging from 128 to 16384. The kernels differ in the multiplication algorithm and the type of optimizations used, ranging from naïve multiplication in Kernel 1 to the most optimized version in Kernel 10. The matrix sizes are chosen to represent a range of computational intensities, from relatively small matrices to very large ones that challenge the capabilities of modern GPUs. We use matrix multiplication as the benchmark since it represents the most fundamental operation in a myriad of GPU workloads. GPUBurn also performs matrix multiplication but with TENSOR cores and hence stresses them. LLM inferencing benchmarks test the computational demands of AI workloads, particularly in scenarios requiring substantial processing power for real-time inference. Together, these benchmarks provide a comprehensive assessment of GPU performance under conditions that closely mimic real-world applications in scientific computing and artificial intelligence.
@@ III-A Analysis of CUDA Workloads
Fig. 1 plots the power distribution for different kernels on V100. (In this set of experiments, clock frequency was set to the maximum possible value of 1380 MHz.) The power consumption of Kernel 1 is concentrated between 50 and 125 W, with the density peaking at $\sim$125 W. For Kernel 2, with power ranging from 100 to 200 W, the density peaks significantly at $\sim$150 W, indicating that this kernel typically consumes more power than Kernel 1. The sharp peak suggests that the power usage is more consistent during this kernel’s operation. Kernel 3’s power range of 100 to 200 units is similar to that of Kernel 2. However, the density distribution shows a broader peak. Kernels 4–7 exhibit similar behavior to Kernel 3 and are hence omitted. Kernels 8–10 show almost similar behavior where power is mostly distributed from 100 to 200 W and we thus use only Kernel 8 to represent all three for brevity. The density peak, however, is slightly lower and broader, suggesting a more distributed power consumption pattern. The subtle variations and the differences in the peaks and spreads of the density curves of these three kernels is an indication of the diversity in the kernel operations that are performed on the variable matrices. Fig. 2 drills down into each kernel’s power by the matrix size. In each scenario, the power increases significantly initially as the matrix size rises, beyond which it saturates. Saturation could be either due to compute or memory becoming a bottleneck or because the GPU’s power limit is reached, at which point the SM frequency is automatically lowered.
Figure 2: Distribution of Power across matrix sizes
@@
We next analyze GPU utilization metrics to account for the differences in the power consumption of kernels. A correlation analysis of about 15 metrics with POWER revealed significant correlation for SMACT, FP32A, and DRAMA. We had locked the GPU clock to 1380 MHz, and hence, there was no correlation between POWER and SMCLK. Since SMACT is a coarse metric that tracks the fraction of time at least one core of any SM is active, it does not provide much insights and hence has been omitted in the analysis of this section.
The FP32A heatmaps of Figs. 3 and 4 indicate the utilization ranges for FP32 CUDA cores for the various kernels, while Figs. 5 and 6 depict the ranges for DRAM usage. FP32A values are higher for larger matrices as expected since they can utilize a larger fraction of SM cores concurrently. It can also be seen that power consumption increases as FP32A increases. The DRAMA heatmaps show that DRAM usage, similar to FP32 activity, increases with matrix size, as larger matrices occupy a larger amount of memory and also lead to larger transfers of data between the memory and on-chip caches. However, unlike FP32A, there is no noticeable trend to larger increase in Kernels 8–10. This is because these kernels are better optimized to eliminate redundant data fetches. In conjunction with the increase in DRAM activity, there is a noticeable rise in power consumption as matrix sizes grow. The heatmaps highlight that kernels with higher DRAM usage also tend to exhibit higher power consumption, indicating a strong correlation between memory activity and energy demands. DRAMA coupled with FP32A highlights the critical role of both memory and processing power in determining overall power consumption.
@@
Figure 6: Power distribution for DRAMA / Kernels
Figure 7: GPU Power vs. FP32A and DRAMA for MATMUL kernels
Utilization vs. Power Characteristics
The 3D plot in Fig. 7 illustrates the interplay between DRAMA, FP32A, and POWER for different GPU kernels. For Kernel 1, a gradual increase in power consumption is observed as both FP32A and DRAMA rise. The values for FP32A are confined to the range of 0.01 to 0.05, with POWER spanning from 60 to 120 W. For Kernel 2, the power levels are noticeably higher, ranging from 100 to 200 W, as FP32A is in the [0.1,0.3] range. DRAMA values remain below 0.15, indicating that this metric contributes less significantly to power variation. Kernel 3 shows a similar trend. For Kernels 8–10, with a broader range for both FP32A (0.0 to 0.8) and DRAMA (0.0 to 0.3), power varies significantly, reaching up to 200 W. A common trend across all the kernels is that power saturates with saturation of FP32A with any further increase in DRAMA (after FP32A saturation) having no perceptible impact. Also, for all the kernels, power increases linearly with FP32A, but with different slopes. The slopes are the highest for Kernels 1–3, which are the least optimized. Note also that there does not seem to be any relation between DRAMA and the slope, since the slope of Kernel 3 with lower DRAMA than Kernels 8–10 is higher, but is lower than Kernel 1 whose DRAMA is lower than itself. This indicates that simple regression models using these and other DCGM metrics are unable to predict power with good accuracy. Building better models is discussed in a later section.
@@
Figure 9: Power distribution for heterogeneous workloads
@@
Our next set of experiments is to determine the impact to power due to concurrent execution of kernels that stress different CUDA cores. To eliminate the influence of memory, we used kernels performing vector operations on FP32 and FP64 data and limited the data sizes so that the data is held on-chip entirely. We first individually exercised the FP32 and FP64 cores, controlling their utilizations via the total number of threads dispatched at a time. Power consumption by core utilization for these independent executions are shown by the blue and red plots, resp., in Fig. 10. The relationship is linear for both due to the absence of DRAM accesses. We next executed the two workloads concurrently, increasing their utilizations systematically. The utilization-power relation for the combined workload is also shown in the same figure (by the green plot). (The total utilization exceeds 1.0, since the utilization for each core type can be up to 1.0.) Note that the combined power for a given total utilization is less than the power consumed by each workload at that utilization level, which by itself suggests that the power consumption of the two workloads may not be additive. For validation, we used linear regression models constructed for the individual workloads (blue and red plots) and used them to compute the expected power for their combined execution. The purple plot shows this expected power, which is much higher than the actual power consumed, further corroborating that FP64 and FP32 powers are not additive. However, this behavior was not so pronounced when both workloads used either FP32 or FP64 cores only. Thus, additivity of workloads cannot be expected to hold at all times, which argues against the the possibility of simple and lightweight power partitioning approaches for MIG.

### v2

@@ Abstract
Efficient power management in cloud data centers is essential for reducing costs, enhancing performance, and minimizing environmental impact. GPUs, critical for tasks like machine learning (ML) and GenAI, are major contributors to power consumption. NVIDIA’s Multi-Instance GPU (MIG) technology improves GPU utilization by enabling isolated partitions with per-partition resource tracking, facilitating GPU sharing by multiple tenants. However, accurately apportioning GPU power consumption among MIG instances, which is required for proper energy and carbon accounting and management, remains challenging due to lack of hardware support. This paper addresses this challenge by developing software methods to estimate power usage per MIG partition. We analyze NVIDIA GPU utilization and power consumption metrics and find that light-weight methods with good accuracy can be difficult to construct. We hence explore the use of ML-based power models to enable accurate, partition-level power estimation. Our findings reveal that a single generic offline power model or modeling method is not applicable across diverse workloads, especially with concurrent MIG usage, and that workload-specific models constructed using partition-level utilization metrics of workloads under execution can significantly improve accuracy. We also propose scaling the MIG power predicted by the models using total measured GPU power, when available, to eliminate aggregate error and reduce error in per-MIG estimations. Using NVIDIA A100 GPUs, we demonstrate and evaluate our approach for accurate partition-level power estimation for workloads including matrix multiplication and Large Language Model inference, contributing to transparent and fair carbon reporting.
@@ I Introduction
Cloud data centers have become the backbone of modern computing, enabling a wide spectrum of applications and services, from social media platforms to complex scientific computations in a cost-efficient manner[1, 2]. As data centers scale, efficient power management is increasingly critical to curb rising energy costs and meet global carbon emission regulations [3, 4]. It also helps reduce operational overhead, improve performance, and lower environmental impact. Among the various components of a data center, Graphics Processing Units (GPUs) play a critical role, particularly in high-performance computing tasks such as machine learning, data analytics, and scientific simulations. With the increasing adoption of GenAI-based applications, the projected growth for data center GPU’s during 2023-28 is 34.6%[5], necessitating the development of efficient strategies for improving their utilization levels and energy efficiency.
@@
We next consider increasingly complex methods for our task at hand. Our first approach is to construct a universal power model that is applicable for a variety of common workloads and use it for independently estimating each partition’s power. It turns out that with ML methods such as Extreme Gradient Boosting (XGBoost)[11], errors can be high with unified models constructed using data sets from dissimilar multiple workloads, in non-MIG, full GPU mode. However, for unified models constructed from similar workloads, the error is substantially reduced. This highlights the potential advantages of applying workload-aware models, which we examine further. Interestingly, we demonstrate that individual workload-aware models do not significantly reduce prediction error when applied to MIG power prediction, whilst a unified model adds to predictivity. Finally, as a measure of improving accuracy, with all the methods, we scale the model-estimated MIG power values using aggregate measured GPU power, when available, and evaluate its efficacy.
@@
  • Highlighting the challenges posed by the problem of partitioning GPU power among MIG instances and showing that lightweight approaches that can be implemented in a low-level module may not be possible.
  • Design and evaluation of ML-based approaches to partition GPU power among MIG instances and a scaling method to lower errors.
@@ II Overview of NVIDIA GPU Architecture
Graphics Processing Units (GPUs), originally intended for accelerating computer graphics[12], have emerged as a cornerstone of high-performance computing due to their highly-parallel processor architecture, consisting of hundreds to thousands of small cores. In NVIDIA GPUs [^1] Unless otherwise specified, GPU henceforth refers to NVIDIA’s GPUs., the processing cores are organized into Streaming Multiprocessors (SMs)[13, 14, 10, 15]. Each SM consists of two types of cores, CUDA[16] and Tensor cores[17, 18], to support arithmetic operations of different types and precisions, such as 16-, 32-, and 64-bit floating points (FP16, FP32, and FP64 resp.) and 8- and 16-bit integers (INT8 and INT16, resp.). CUDA cores handle scalar operations per cycle, while Tensor cores accelerate AI workloads with mixed-precision matrix operations. Due to their simpler design, CUDA cores far outnumber Tensor cores in an SM. GPUs feature sophisticated L1/L2 caches, shared memory, and high-bandwidth off-chip memory, enhancing performance in data-intensive tasks.
The above hardware support for parallelism is accompanied by a matching parallel software programming and execution model based on warps. A warp, typically a group of 32 threads, executes the same instruction on different data within an SM, which schedules and manages their execution. If a warp stalls, the SM switches to a ready warp, minimizing latency and keeping cores fully utilized. This fine-grained multithreading, following the SIMT/SIMD model, maximizes throughput by exploiting data-level parallelism in workloads like AI.
GPU performance is assessed through hardware metrics such as core clock speed (SMCLK), memory utilization (DRAMA), CUDA core (FP64A, FP32A, FP16A, INT16A) and tensor core (TENSO) usage, and power consumption (POWER) [19]. These metrics help optimize performance and energy efficiency. In data centers, they can be monitored with minimal overhead using NVIDIA’s Data Center GPU Manager (DCGM) [20].
NVIDIA Nsight Compute[21] offers detailed, kernel-level GPU profiling, analyzing memory access, warp execution, and thread divergence. While valuable for workload optimization, it incurs higher overhead and is suited for development rather than real-time monitoring. This study evaluates both DCGM and Nsight metrics for power prediction, balancing detail and efficiency for cloud data centers.
@@
Multi-Instance GPU (MIG): Modern data center GPUs offer up to a thousand TFLOPS[22], but workloads may not be capable of fully utilizing this capacity. To improve the overall utilization, extend access to wider set of users, and also reduce cost, different sharing modes namely, time-slicing, multi-process service (MPS)[23], and multi-instance GPU (MIG)[7] are supported by NVIDIA, with only MIG ensuring complete performance and memory isolation. MIG divides SMs into seven and memory into eight slices, which can be combined into configurable GPU instances or partitions. The configurations, referred to as MIG profiles, permissible for the MIG partitions in an A100 with 80 GB memory are provided in Table I. As discussed in Sec. I, in MIG mode, GPU utilization metrics are reported at the partition level, while power consumption is available only for the full GPU. The rest of the paper discusses methods for distributing the aggregate power among the partitions.
@@ III GPU Power-Performance Characterization Studies and Modeling
In this section, we examine the various factors that affect power consumption in GPUs using NVIDIA A100 and V100 systems. We use different workloads such as various CUDA matrix multiplication kernels (MATMUL)[24], a TENSOR workload based on GPU Burn benchmark (GPUBurn)[25], and Large Language Model (LLM) inference systems[26]. MATMUL evaluates ten matrix multiplication kernels (Kernels 1–10) across matrix sizes from 128 to 16,384, varying in algorithm and optimization levels, ranging from naïve multiplication in Kernel 1 to the most optimized version in Kernel 10. These sizes capture diverse computational demands, making matrix multiplication an ideal benchmark for GPU workloads. GPUBurn stresses Tensor cores with matrix multiplication, while LLM benchmarks assess AI inference workloads. Together, they provide a comprehensive GPU performance evaluation for scientific computing and AI applications.
@@ III-A Analysis of CUDA Workloads
Figure 2: Distribution of Power across matrix sizes
@@
Fig. 1 plots the power distribution for different kernels on V100 by setting clock frequency to its maximum value 1380 MHz. The power consumption of Kernel 1 is concentrated between 50 and 125 W, with the density peaking at $\sim$125 W. For Kernel 2, with power ranging from 100 to 200 W, the density peaks significantly at $\sim$150 W, indicating higher power consumption than Kernel 1. The sharp peak suggests that the power usage is more consistent during this kernel’s operation. Kernel 3’s power range of 100 to 200 units is similar to that of Kernel 2 barring a broader peak. Kernels 4–7 exhibit similar behavior to Kernel 3 and are hence omitted. Kernels 8–10 also share a 100–200 W distribution, with Kernel 8 only depicted for brevity. The density peak, however, is slightly lower and broader, suggesting a more distributed power consumption pattern. The subtle variations and the differences in the peaks and spreads of the density curves of these three kernels reflect variations in kernel operations on different matrix sizes. Fig. 2 examines power by matrix size with power initially rising with matrix size before saturating, likely due to compute or memory bottlenecks or GPU power limits, which trigger automatic SM frequency scaling.
@@
We next analyze GPU utilization metrics to account for the differences in the power consumption of kernels. A correlation analysis of about 15 metrics with POWER revealed significant correlation for SMACT, FP32A, and DRAMA. We had locked the GPU clock to 1380 MHz, and hence, there was no correlation between POWER and SMCLK. Since SMACT is a coarse metric that tracks the fraction of time at least one core of any SM is active, it does not provide much insights and hence has been omitted in the analysis of this section. The FP32A heatmaps of Fig. 3 indicate the utilization ranges for FP32 CUDA cores for the various kernels, while Fig. 4 depict the ranges for DRAM usage. FP32A values are higher for larger matrices since they can utilize a larger fraction of SM cores concurrently. It can also be seen that power consumption increases with FP32A.
The DRAMA heatmaps show that DRAM usage, similar to FP32 activity, increases with matrix size, as larger matrices occupy a larger amount of memory and also lead to larger transfers of data between the memory and on-chip caches. In conjunction with the increase in DRAM activity, there is a noticeable rise in power consumption as matrix sizes grow. The heatmaps highlight that kernels with higher DRAM usage also tend to exhibit higher power consumption, indicating a strong correlation between memory activity and energy demands. DRAMA coupled with FP32A highlights the critical role of both memory and processing power in determining overall power consumption.
@@
Figure 6: GPU Power vs. FP32A and DRAMA for MATMUL kernels
Utilization vs. Power Characteristics
The 3D plot in Fig. 6 illustrates the interplay between DRAMA, FP32A, and POWER for different GPU kernels. For Kernel 1, a gradual increase in power consumption is observed as both FP32A and DRAMA rise. The values for FP32A are confined to the range of 0.01 to 0.05, with POWER spanning from 60 to 120 W. For Kernel 2, the power levels are noticeably higher, ranging from 100 to 200 W. DRAMA values remain below 0.15, indicating that this metric contributes less significantly to power variation. Kernel 3 shows a similar trend. For Kernels 8–10, with a broader range for both FP32A (0.0 to 0.8) and DRAMA (0.0 to 0.3), power varies significantly, reaching up to 200 W. Across all kernels, power saturates after peaking with FP32A, with further increases in DRAMA having no impact. Power scales linearly with FP32A but with varying slopes—highest for Kernels 1–3, which are the least optimized. Note also that there does not seem to be any relation between DRAMA and the slope, since the slope of Kernel 3 with lower DRAMA than Kernels 8–10 is higher, but is lower than Kernel 1 whose DRAMA is lower than itself.
@@
We now analyze the power impact of concurrently executing FP32 and FP64 vector kernels while minimizing memory influence by keeping data entirely on-chip. By independently exercising FP32 and FP64 cores and varying thread dispatch, we control utilization to assess their individual and combined power consumption effects. Power consumption by core utilization for these independent executions are shown by the blue and red plots, resp., in Fig. 7. The relationship is linear for both due to the absence of DRAM accesses. We next executed the two workloads concurrently, increasing their utilizations systematically. The utilization-power relation for the combined workload is also shown in the same figure (by the green plot). (The total utilization exceeds 1.0, since the utilization for each core type can be up to 1.0.)
@@
Note that the combined power for a given total utilization is less than the power consumed by each workload at that utilization level, which by itself suggests that the power consumption of the two workloads may not be additive. For validation, we used linear regression models constructed for the individual workloads (blue and red plots) and used them to compute the expected power for their combined execution. The purple plot shows this expected power, which is much higher than the actual power consumed, further corroborating that FP64 and FP32 powers are not additive. However, this behavior was not so pronounced when both workloads used either FP32 or FP64 cores only. Thus, additivity of workloads cannot be expected to hold at all times, which argues against the the possibility of simple and lightweight power partitioning approaches for MIG.